OpenStudy (anonymous):

Find the mean. 5.39, 8.35, 3.81 4.54

OpenStudy (anonymous):

(5.39+8.35+3.81+4.54)/4=22.09/4=5.5225

OpenStudy (texaschic101):

The mean is the average. It is found by adding up all the numbers, and then dividing by the total number that you have